OpenStudy (anonymous):

Annie wrote the equation shown below. 3(3y) = (3 • 3) y Which property did Annie apply? associative property of multiplication commutative property of multiplication multiplication property of equality distributive property of multiplication

hero (hero):

Which do you believe is the correct answer?

OpenStudy (anonymous):

i chose the last one

hero (hero):

Distributive Property is a(b+c) = ab + bc That's not exactly what we have here.

hero (hero):

There's only one property that switches parentheses like that

OpenStudy (anonymous):

hmm ok so maybe its associative prop of multiplication

hero (hero):

Yes, as you know, associative property is a(bc) = (ab)c In this case we have 3(3y) = (3*3)y a direct correspondence of the property

OpenStudy (anonymous):

ok thanks do you mind helping me with just a feew more Q: if it not much to ask for

hero (hero):

I'm kinda busy with my own work. Maybe another time. Nonetheless, you can tag or ping any other user you feel might be able to help. Another thing people do to get attention for their questions is copy and paste the link to their question in the math chat.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well i dont want to hold you back then